OpenStudy (anonymous):

alfred deposited $893.78 in a savings account that earns 2.2% simple interest. What is alfreds account balance after seven years?

OpenStudy (anonymous):

@terenzreignz

OpenStudy (mathstudent55):

The formula for simple interest is: I = prt I = interest earned p = principal amount (amount invested) r = interest rate written as a decimal t = time in years Calculate the interest using the formula above and add to the principal ($893.78).

OpenStudy (anonymous):

What is my formula ?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

I have no idea ? Is it right ?

OpenStudy (anonymous):

isnt the answer 1031.42? 893.78*.022=19.6632(7)+893.78=1031.42 you multiply your base amount (893.78) by the interest in decimal form (.022) whih equals 19.66316 (this is the interest over only 1 year..so you have to multiply it over the course of 7 years which is 137.642) then add it to the starting base amount 893.78..so the answer is 1031.42

OpenStudy (mathstudent55):

I = prt I = $893.78 * 0.022 * 7 I = $137.64 $893.78 + $137.64 = $1031.42